OpenStudy (anonymous):

what mass, in grams, of No can form from 7.108 g of nh3

OpenStudy (ehsan18):

Its probably only N, No does not symbolize any element, Molar Mass of Ammonia(NH3)=14+1(3)=17 Given Mass=7.108 Mass of nitrogen in 17g of Ammonia = 14g Mass of nitrogen in 1g of ammonia = 14/17 Mass of nitrogen in 7.108g of ammonia=14/17 * 7.108 Hence the answer is 5.85

OpenStudy (jfraser):

except nitrogen gas is \(N_2\), not just \(N\). You need to know what the balanced reaction is before you can do any kind of stoichiometric calculations

OpenStudy (ehsan18):

@JFraser is the solution posted by me wrong?

OpenStudy (jfraser):

probably, since you don't know the balanced reaction. Maybe the stoichiometric ratio is 1:1, but maybe it's not. This question is impossible to answer properly without more information

OpenStudy (cuanchi):

If we assume that "No" is nitrogen. All the Nitrogen doesn't matter if is molecular or atomic (the total mass is the same) is coming from the NH3, there is not other source of nitrogen, we don't need stoichiometry to solve this problem and I think that the approach from @Ehsan18 is correct. It is a percent composition kind of problem more than a stoichiometry problem. If "No" is "NO" (nitrogen oxide), we will need some adjustment in the calculations.